# Baorui Ma,  Huachen Gao,  Haoge Deng,  Zhengxiong Luo,  Tiejun Huang,  Lulu Tang,  Xinlong Wang

### v1

@@ 1 Introduction
MVD models have recently gained widespread attention due to their advantages of integrating the generative capabilities of 2D diffusion models while maintaining consistency across multiple views [74, 50]. Typically, these models rely on precise camera poses [74, 21, 30, 48, 60] or warped images rendered according to camera position [112, 87] as conditional inputs to guide 3D-consistent view generation. We refer to these conditions, derived from pose or 3D annotations, as 3D-inductive conditions. However, annotating web-scale videos is prohibitively costly, or even intractable in some cases, posing significant challenges for scaling. To address this, we propose a novel, pose-free visual-condition derived from pixel-space hints within videos. It is a purely 2D-inductive visual signal, created by introducing time-dependent noise to masked input videos, free from any 3D-inductive bias. This enables training MVD model at scale, without requiring pose annotations.
@@ 3.2 Visual Conditional Multi-View Diffusion Model
Principle of Visual-condition. A desirable visual-condition should meet the following criteria: a) it can be constructed without the need for additional 3D annotations, b) it is independent of specific downstream tasks, and c) it offers sufficient generalization to support various task-specific visual conditions, enabling precise control of camera movements.
@@
Ideally, this visual-condition can be derived from pixel-space hints within the original videos, implicitly guiding the model to learn camera control. Moreover, it should be robust enough to handle domain gaps between task-specific visual cues and pixels extracted from video data. For example, in warping-based generation, warped images often suffer from issues like self-occlusions, artifacts, and distortions, creating a significant gap compared to real video data as shown in Fig.4 and Fig.5. Based on this, we propose constructing the visual-condition by applying masks and noise to the input video data. Our design principles are:
@@
  • Masking: Corrupting video data through random masking reduces reliance on direct pixel-space visual signals, helping the model partially mitigate the domain gap between task-specific visual cues and video data.
  • Noise Addition: Adding noise to video data to approximate a Gaussian distribution. When applied to different downstream tasks, task-specific visual inputs are similarly noised, aligning their distribution with a Gaussian profile, and further bridging the gap between video data and task-specific visual input.
Time-dependent Visual Condition. We initially add noise and random irregular masks to video data, creating what we call ‘corrupted’ video data $C_{t}$, as defined in Eq.2. This corrupted data is then used as the visual-condition for MVD model training. A key challenge of this strategy lies in determining the appropriate level of perturbation. If too much noise is added, the conditional signals become ineffective, resulting in poor visual quality and inaccurate camera control. Conversely, if insufficient noise is added, the corrupted data retains substantial details from the target images, causing the model to overly rely on visual hints from video data.
To overcome this, we further introduce a time-dependent noise distribution applied to video data. The core idea is to apply higher noise levels at larger time steps, effectively disrupting video data to prevent over-reliance on it. As the time step decreases, the noise level is reduced, providing cleaner conditional signals to facilitate content generation. However, as the noise level decreases, the risk of information leakage increases, causing the distribution to deviate from a Gaussian distribution. To mitigate this, we propose gradually replacing corrupted data $C_{t}$ with noisy latent $X_{t}$ as the visual-condition (see Eq.3) , encouraging the model to primarily rely on $X_{t}$ for generation while minimizing dependence on pixel-space signals from video data as timestep decreases. Additionally, $X_{t}$ is predicted by the model, thus exhibiting task-agnostic property, as formalized by:
$C_{t}=\sqrt{\bar{\alpha}_{t^{\prime}}}(1-M)\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ (2)
@@
where $M=\left\{m^{0:S}\cup m^{S+1:N}\right\}$, with $m^{0:S}$ as the zero matrix, keeping the reference images $Y_{0}$ unmasked, and $m^{S+1:N}$ as random irregular masks applied to the target images $G$. $t^{\prime}=f(t)$, function $f$ is strictly monotonically increasing and requires $t^{\prime}<t$, which guarantees that $C_{t}$ contains at least as much information as $X_{t}$ at earlier timesteps. Here, $\bar{\alpha}_{t^{\prime}}$ are variances used in DDIM [78], and $V_{t}=\left\{v_{t}^{i}\right\}_{i=1}^{N}$ represents our final visual-condition, which is a mixture of $C_{t}$ and $X_{t}$, concatenated with masks $M$ along the channel dimension. The term $W_{t}$ denotes a balancing weight that decreases monotonically with timestep $t$, ranging from 1 to 0. In practice, an additional processing step assigns $v_{t}^{0:S}$ to the reference images $Y_{0}$ directly, in order to inject the clean information of $Y_{0}$ into the model, facilitating alignment between the predicted images and the reference images. Consequently, Eq.1 can be reformulated as $\mathbb{E}_{X_{0},Y_{0},\epsilon,t}\left[\left\lVert\epsilon_{\theta}(X_{t},Y_{0},V_{t},t)-\epsilon\right\rVert_{2}^{2}\right]$. More detailed definitions of $f(t)$ and $W_{t}$ are provided in the Appendix C.3.
Model Architecture. Our model architecture is based on video diffusion model [5]. However, we removed the time embedding, as we aim for the model to control the camera movement purely through visual conditions, rather than inferring movement trends based on temporal cues. To further minimize the effect of temporality, we shuffle the frames in each video clip, treating the data as unordered $X_{0}$. Specifically, we randomly select a subset of frames from a video clip as reference images, with the remaining frames as target images. The number of reference images is randomly selected to accommodate different downstream tasks. The multi-view diffusion model is optimized by calculating the loss only on the target images, as described in Eq.1. Additional details on the model architecture, including the design of self-attention layers, Zero-Initialize, trainable parameters, noise schedule, and cross-attention, are provided in the Appendix C.1.
@@ 3.3 Visual Conditional 3D Generation
$\bm{W}_{u,v}=diag(w_{1},w_{2},...,w_{m}),w_{i}=\frac{1}{\sqrt{2\pi}}\exp(-\frac{\mathrm{dist}_{i}^{2}}{2b^{2}}),$ (5)
@@ 4.1 Single View to 3D
Results. The quantitative comparison results are presented in the top rows of Tab. 2. Only average metrics for the easy and hard sets are reported here, detailed values are available in the Appendix D.1. The results for ViewCrafter* are comparable to those reported in its original paper, confirming successful alignment between our method and the baselines. Numerically, our approach outperforms all baseline methods across all metrics. Specifically, compared to the re-implemented ViewCrafter, our approach achieves a 4.63 dB improvement, demonstrating its capability to generate high-quality novel views. PSNR further demonstrates significant gains, indicating our proposed visual-condition enables precise camera control. Qualitative results are shown in the top rows of Fig. 4. See3D generates high-quality, realistic content within minutes. Dispite limited visual cues provided by the warped images, our method produces more reliable and realistic results with fewer artifacts.
@@ C.1 Model Architecture
The main backbone of See3D model is based on the structure of 2D diffusion models but integrates 3D self-attention to connect the latents of multiple images, as shown in prior work [74]. Specifically, we adapt the existing 2D self-attention layers of the original 2D diffusion model into 3D self-attention by inflating different views within the self-attention layers. To incorporate visual conditions, we introduce the necessary convolutional kernels and biases using Zero-Initialize [76].
@@
The model is initialized from a pretrained 2D diffusion model [65] and fine-tuned with all parameters, leveraging FlashAttention for acceleration. In accordance with prior work [73], switching from a scaled-linear noise schedule to a linear schedule is essential for achieving improved global consistency across multiple views. Additionally, we implement cross-attention between the latents of multiple views and per-token CLIP embeddings of reference images using a linear guidance mechanism [79]. For training, we randomly select a subset of frames from a video clip as reference images, with the remaining frames serving as target images. The number of reference images is randomly chosen to accommodate different downstream tasks. The multi-view diffusion model is optimized by calculating the loss only on the target images, as outlined in Eq. 1.
C.2 Training Configurations
We initialize the See3D model from MVDream [74] and employ a progressive training strategy. First, the model is trained at a resolution of 512 $\times$ 512 with a sequence length of 5. This phase involves 120,000 iterations, using 1 reference view and 4 target views. Due to the relatively small sequence length, a larger batch size of 560 is used to enhance stability and accelerate convergence. Next, the sequence length is increased to 16, and the model is trained for 200,000 iterations with 1 or 3 reference views and 15 or 13 target views, maintaining the resolution of 512 $\times$ 512. In this phase, the batch size is reduced to 228. Finally, a multi-view super-resolution model is trained using the same network structure. It takes the multi-view predictions from See3D as input and outputs target images with multi-view consistency at a resolution of 1024 $\times$ 1024, using a batch size of 114. In all stages, all parameters of the diffusion model are fine-tuned with a learning rate of 1e-5.
@@ C.2 Training Configurations
Additionally, we render some multi-views or extract clips from datasets such as Objaverse [14], CO3D [69], RealEstate10k [117] , MVImgNet [113], and DL3DV [45] datasets, forming a supplemental 3D dataset with fewer than 0.5M samples, please refer to Section E.2 for details on analysis and ablation. During training, this supplemental data is randomly sampled and incorporated into our WebVi3D dataset ($\sim$16M). To enhance training efficiency, we utilize FlashAttention [13] alongside DeepSpeed with ZeRO stage-2 optimizer [68] and bf16 precision. We also implement classifier-free guidance (CFG) [27] by randomly dropping visual conditions with a probability of 0.1. The See3D model is trained on 114 $\times$ NVIDIA-A100-SXM4-40GB GPUs over approximately 25 days using a progressive training scheme. During inference, a DDIM sampler [78] with classifier-free guidance is employed.

### v2

@@ 1 Introduction
MVD models have recently gained widespread attention due to their advantages of integrating the generative capabilities of 2D diffusion models while maintaining consistency across multiple views [80, 56, 51, 88, 128]. Typically, these models rely on precise camera poses [80, 23, 33, 54, 66, 2, 110, 111, 28, 124, 45] or warped images rendered according to camera position [121, 95] as conditional inputs to guide 3D-consistent view generation. We refer to these conditions, derived from pose or 3D annotations, as 3D-inductive conditions. However, annotating web-scale videos is prohibitively costly, or even intractable in some cases, posing significant challenges for scaling. To address this, we propose a novel, pose-free visual-condition derived from pixel-space hints within videos. It is a purely 2D-inductive visual signal, created by introducing time-dependent noise to masked input videos, free from any 3D-inductive bias. This enables training MVD model at scale, without requiring pose annotations.
@@ 3.2.1 Principle of Visual-Condition
A desirable visual-condition should meet the following criteria: a) it can be constructed without the need for additional 3D annotations, b) it is independent of specific downstream tasks, and c) it offers sufficient generalization to support various task-specific visual conditions, enabling precise control of camera movements.
@@
Ideally, this visual-condition can be derived from pixel-space hints within the original videos, implicitly guiding the model to learn camera control. Moreover, it should be robust enough to handle domain gaps between task-specific visual cues and pixels extracted from video data. For example, in warping-based generation, warped images often suffer from issues like self-occlusions, artifacts, and distortions, creating a significant gap compared to real video data as shown in Fig.4 and Fig.5.
3.2.2 Time-dependent Visual Condition
Building on the analysis above, we propose constructing the visual-condition by applying masks, noise, and mixture to the input video data.
@@ 3.2.2 Time-dependent Visual Condition
Random Masking: We first corrupt target images $G$ through random irregular masking to reduce reliance on direct pixel-space visual signals, helping the model partially mitigate the domain gap between task-specific visual cues and video data. Meanwhile, we keep the reference images $Y_{0}$ clean to provide effective appearance signals.
Time-dependent Noise: We further add noise to video data to approximate a Gaussian distribution. For downstream tasks, task-specific visual inputs are similarly noised, aligning their distributions with this Gaussian profile and further bridging the gap between video data and task-specific inputs. A key challenge lies in determining the optimal noise level: excessive noise weakens conditional signals, resulting in poor visual quality and inaccurate camera control, whereas insufficient noise preserves too many details from the target images, causing the model to over-rely on visual hints from the video data.
Previous studies [66, 83, 34, 127, 18] have explored modulating noise levels by adding noise to input data. Notably, as pointed out in the previous work [127], diffusion models tend to over-rely on the conditional image at larger time steps, leading to signal leakage. Inspired by this[127], we introduce time-dependent noise to the corrupted target images. In addition, we develop a function $t^{\prime}=f(t)$ to regulate signal leakage, preventing excessive noise from completely obscuring visual cues and disabling camera control. Specifically, we define:
$C_{t}=\sqrt{\bar{\alpha}_{t^{\prime}}}(1-M)\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}\bm{\epsilon}.\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (2)
Here, $f$ is a strictly monotonically increasing function, ensuring $t^{\prime}<t$, so that $C_{t}$ contains at least as much information as $X_{t}$ at earlier timesteps. $\bar{\alpha}_{t^{\prime}}$ are the variances used in DDIM [85]. A detailed explanation of $f(t)$ can be found in Appendix C.3.
Time-dependent Mixture: However, as $t$ decreases, lower noise levels increase the risk of signal leakage, causing a domain gap between the video data and task-specific visual condition distributions. To address this issue, we propose gradually replacing the corrupted data $C_{t}$ with noisy latent variables $X_{t}$ as timestep decreases. This encourages the model to rely more on pixel-space signals from video data at larger time steps, and transition to $X_{t}$ at smaller timesteps. To achieve this, we further introduce a weighting factor $W_{t}\in[0,1]$, which decreases monotonically with the timestep $t$, to combine $C_{t}$ and $X_{t}$. Formally, our final visual-condition is defined as:
@@
where $M=\left\{m^{0:S}\cup m^{S+1:N}\right\}$, with $m^{0:S}$ as a zero matrix, keeping the reference images $Y_{0}$ unmasked, and $m^{S+1:N}$ as random irregular masks applied to the target images $G$. $V_{t}=\left\{v_{t}^{i}\right\}_{i=1}^{N}$ represents a mixture of $C_{t}$ and $X_{t}$, concatenated with masks $M$ along the channel dimension. In practice, an additional processing step assigns $v_{t}^{0:S}$ to the reference images $Y_{0}$ directly, in order to inject the clean information of $Y_{0}$ into the model, facilitating alignment between the predicted images and the reference images. Consequently, Eq.1 can be reformulated as $\mathbb{E}_{X_{0},Y_{0},\epsilon,t}\left[\left\lVert\epsilon_{\theta}(X_{t},Y_{0},V_{t},t)-\epsilon\right\rVert_{2}^{2}\right]$. A more detailed definition of $W_{t}$ is provided in Appendix C.3.
3.2.3 Model Architecture
Our model architecture is based on video diffusion model [6]. However, we removed the time embedding, as we aim for the model to control the camera movement purely through visual conditions, rather than inferring movement trends based on temporal cues. To further minimize the effect of temporality, we shuffle the frames in each video clip, treating the data as unordered $X_{0}$. Specifically, we randomly select a subset of frames from a video clip as reference images, with the remaining frames as target images. The number of reference images is randomly selected to accommodate different downstream tasks. The multi-view diffusion model is optimized by calculating the loss only on the target images, as described in Eq.1. Additional details regarding the model architecture, including the design of self-attention layers, Zero-Initialize, trainable parameters, noise schedule, and cross-attention, can be found in the Appendix C.1.
@@ 3.3 Visual Conditional 3D Generation
$\begin{gathered}\bm{W}_{u,v}=diag(w_{1},w_{2},...,w_{m}),\\ w_{i}=\frac{1}{\sqrt{2\pi}}\exp(-\frac{\mathrm{dist}_{i}^{2}}{2b^{2}}),\end{gathered}$ (5)
@@ 4.1 Single View to 3D
Results. The quantitative comparison results are presented in the top rows of Tab. 2. Only average metrics for the easy and hard sets are reported here, detailed values are available in the Appendix D.1. The results for ViewCrafter* are comparable to those reported in its original paper, confirming successful alignment between our method and the baselines. Numerically, our approach outperforms all baseline methods across all metrics. Specifically, compared to the re-implemented ViewCrafter, our approach achieves a 4.63 dB improvement, demonstrating its capability to generate high-quality novel views. PSNR further demonstrates significant gains, indicating our proposed visual-condition enables precise camera control. Qualitative results are shown in the top rows of Fig. 4. See3D generates high-quality, realistic content within minutes. Despite limited visual cues provided by the warped images, our method produces more reliable and realistic results with fewer artifacts.
@@ C.1 Model Architecture
The main backbone of See3D model is based on the structure of 2D diffusion models but integrates 3D self-attention to connect the latents of multiple images, as shown in prior work [80]. Specifically, we adapt the existing 2D self-attention layers of the original 2D diffusion model into 3D self-attention by inflating different views within the self-attention layers. To incorporate visual conditions, we introduce the necessary convolutional kernels and biases using Zero-Initialize [82]. The model is initialized from a pretrained 2D diffusion model [71] and fine-tuned with all parameters, leveraging FlashAttention for acceleration. In accordance with prior work [79], switching from a scaled-linear noise schedule to a linear schedule is essential for achieving improved global consistency across multiple views. Additionally, we implement cross-attention between the latents of multiple views and per-token CLIP embeddings of reference images using a linear guidance mechanism [86]. For training, we randomly select a subset of frames from a video clip as reference images, with the remaining frames serving as target images. The number of reference images is randomly chosen to accommodate different downstream tasks. The multi-view diffusion model is optimized by calculating the loss only on the target images, as outlined in Eq. 1.
C.2 Training Details
Brightness Control. We observe that the visual-condition effectively guides camera movement but cannot control brightness changes, posing a significant limitation. Determining the light source position is particularly challenging with limited observations from single or sparse views. In our real-world test data, camera movement often causes random highlighting or darkening in some regions of scenes, which has a significant impact on pixel-level metrics like PSNR. This issue highlights a key problem: the inability to control brightness undermines the reliability of pixel-level metrics, as brightness variations affect these metrics more than the actual quality of the generated content. To achieve illumination control, 1) we preprocess the training data by converting corrupted images into HSV format, which represents hue, saturation, and brightness. 2) We define a $w\times h$ window and calculate the average brightness difference within this window between the ground truth image and the corrupted data. Using this difference, we apply a scaling factor to the brightness channel of the corrupted data while preserving hue and saturation, before converting the image back to RGB. This ensures brightness adjustment in the visual-condition without leaking color or content from the ground truth.
During training, we randomly drop this preprocessing with a probability of 0.5, enabling the model to infer lighting changes on its own during inference when brightness control is not required. In our evaluation experiments, brightness scaling is applied to the unmasked regions of warped images to align with ground truth, reducing the impact of brightness, and thus yielding a higher correlation between the generated content and pixel-level metrics. Meanwhile, keeping hue and saturation unchanged to avoid content or color leakage. Additionally, the model enables user-controlled brightness adjustments for specific regions in multi-view generation by modifying the visual-condition as needed.
@@ C.2 Training Details
Training Configuration. We initialize the See3D model from MVDream [80] and employ a progressive training strategy. First, the model is trained at a resolution of 512 $\times$ 512 with a sequence length of 5. This phase involves 120,000 iterations, using 1 reference view and 4 target views. Due to the relatively small sequence length, a larger batch size of 560 is used to enhance stability and accelerate convergence. Next, the sequence length is increased to 16, and the model is trained for 200,000 iterations with 1 or 3 reference views and 15 or 13 target views, maintaining the resolution of 512 $\times$ 512. In this phase, the batch size is reduced to 228. Finally, a multi-view super-resolution model is trained using the same network structure. It takes the multi-view predictions from See3D as input and outputs target images with multi-view consistency at a resolution of 1024 $\times$ 1024, using a batch size of 114. In all stages, all parameters of the diffusion model are fine-tuned with a learning rate of 1e-5. Additionally, we render some multi-views or extract clips from datasets such as Objaverse [15], CO3D [75], RealEstate10k [129] , MVImgNet [122], and DL3DV [50] datasets, forming a supplemental 3D dataset with fewer than 0.5M samples, please refer to Section E.2 for details on analysis and ablation. During training, this supplemental data is randomly sampled and incorporated into our WebVi3D dataset ($\sim$16M). To enhance training efficiency, we utilize FlashAttention [14] alongside DeepSpeed with ZeRO stage-2 optimizer [74] and bf16 precision. We also implement classifier-free guidance (CFG) [30] by randomly dropping visual conditions with a probability of 0.1. The See3D model is trained on 114 $\times$ NVIDIA-A100-SXM4-40GB GPUs over approximately 25 days using a progressive training scheme. During inference, a DDIM sampler [85] with classifier-free guidance is employed.